# Blind Image Deblurring using FFT-ReLU with Deep Learning Pipeline Integration

### v1

@@ 13       $scale\leftarrow scale+1$
This blind deconvolution algorithm estimates the latent image $I$ and the blur kernel $k$, which is necessary for obtaining the latent image. Initially, the kernel $k$ is initialized to the coarsest resolution (line 1). Line 1 determines the scale according to the current kernel size. The algorithm iteratively increases the scale. For each scale within a loop, the blurry image $B$ is down-sampled to current scale in line 1. We iterate a predetermined number of times within this setting to refine estimation of $I$ using Equation (6) in line 1, and $k$ using Equation (5) in line 1. Isolated noises are removed by deleting connected components within the kernel that do not exceed a specified threshold (line 1). The kernel is centered after the noise removal in line 1. kernel is then up-sampled with the current scale in line 1. All down-sampling and up-sampling are done using bi-linear interpolation.
@@ 9 $I\_result\leftarrow I_{1}-$ bilateral_filter($\mathrm{diff}$)
For non-blind deconvolution, we estimate an image $I_{1}$ with such priors in lines 2 and 2. Then, we use the L0 norm of the gradient from our Equation (4) to estimate another latent image $I_{2}$ in line 2, since gradient-based methods have been shown to provide accurate results in suppressing ringing artifacts. Similar to ringing suppression methods discussed by [36], we compute a difference map between these two estimated images in line 2. We use bilateral filtering on the computed difference map and subtract the result from $I_{1}$ in line 2, thereby smoothing out artifacts and completing non-blind deconvolution.
4 Empirical Evaluation
This section discusses the parameter settings and provides quantitative results on datasets referenced from [16, 21]. Additionally, it presents visual results across various domains, including natural scenes, low-light conditions, text, and facial images. We also integrate our method as an add-on within a transformer-based deep learning framework and conduct comparative analyses.
4.1 Empirical Settings
@@
We compare our results with both statistical deblurring algorithms (from [36, 11, 17, 49, 32, 10, 44, 13]) and deep-learning based methods (from [41, 5, 18, 19, 53, 51, 52]), where our algorithm was used as a post-processing unit. The metrics chosen for the comparison include PSNR, SSIM, error ratio and inference time. In our experiments, the parameters $\lambda,\mu$ and $\alpha$ are assigned values $0.0043\times 10^{-4},$ 0.004 and $2$, respectively. We compared the convergence of the algorithm for various values of these parameters with respect to kernel similarity, and a more detailed analysis of this is provided in the supplementary material. The value of max_iter is set to be 5 as a trade-off between timing and precision. Our implementation for blind image deblurring on single images is done using Python. [^1] The source code is available on the author’s GitHub pages. Previous blind image deblurring work was mainly done in MATLAB, so translating it allows for GPU access and integration with deep learning techniques.
4.2 Benchmarking with Statistical Deblurring Algorithms
@@
In this section, we evaluate our proposed algorithm with state-of-the-art statistical deblurring algorithms based on the metrics of PSNR, SSIM and error ratio, followed by a visual comparison with those algorithms.
4.2.1 Quantitative Results
We compare our results with others using images from [16] and [21], employing PSNR, SSIM, error ratio and inference time as metrics.
@@ 4.2.1 Quantitative Results
PSNR on Köhler et al. dataset: We computed the average PSNR of our algorithm’s output across 32 images. For each of the 4 images and 8 kernels, there are 199 corresponding ground truth images used for PSNR calculation in [16], against other statistical deblurring algorithms.
Figure 4: Left: PSNR comparison with other algorithms, Right: PSNR with and without our prior
We observe competitive performance when compared with the aforementioned algorithms, and the effect of the presence of our ReLU Sparsity Prior is also conspicuous in Fig. 4.
@@
PSNR SSIM Error ratio Average inference time
[10] 25.63 0.7907 2.6688 1.3951 seconds
[32] 27.54 0.8626 1.2076 109.6088 seconds
[33] 28.38 0.925 15.0949 seconds
[43] 32.445 0.9344 18.61 seconds
[6] 33.145 65.2 seconds
Ours 28.34 0.887 0.3475 8.057 seconds
@@
In Table 1, we show that our method achieves the best error ratio, while achieving a reasonable tradeoff between inference time and PSNR/SSIM.
Inference Time Comparison: We compare our work with [32] and [33] in this section, using datasets from [21, 16, 39]. The images in this datasets are of dimensions $255\times 255$, $800\times 800$ and $1024\times 800$, respectively.
@@
Levin et al. Köhler et al. Sun et al.
[32] 109.609 1550.932 2280.716
[33] 15.095 210.981 269.872
[43] 18.61 85.55
[6] 65.2 755.43
Ours 8.057 74.4301 88.088
Table 2: Inference Times of Algorithms on Different Datasets
@@
Our approach achieves significant speed gains: up to 3 times compared to [33], as presented in Table 2.
4.2.2 Visual Results
In this section, we present a visual comparison with other statistical deblurring algorithms using images from different domains, such as natural images, text images, low-light images, and images containing human face.
@@ 4.2.2 Visual Results
Figure 5: Left to right: input image, followed by results from [10, 45, 32] and our algorithm
@@
Figure 6: Left to right: input image, followed by results from [10, 58, 31] and our algorithm
Figure 7: Left to right: input image, followed by results from [45, 31, 32] and our algorithm
Figure 8: Left to right: input image, followed by results from [9, 10, 31] and our algorithm
Figure 9: Left to right: input image, followed by results from [45, 14, 32] and our algorithm
Figure 10: Left to right: input image, followed by results from [45, 44, 31] and our algorithm
Figure 11: Left to right: input image, followed by results from [45, 30, 32] and our algorithm
Figure 12: Left to right: input image, followed by results from [45, 32, 31] and our algorithm
Regardless of the domain that the image comes from, we obtain consistent results with improved image clarity (as seen in Figures 8, 8, 8, 8, 12, 12, 12, 12), which is a proof of the higher generalisation capabilities of our algorithm compared to other domain-specific methods.
4.3 Benchmarking with Transformer Based Deep Learning Algorithms
We conducted experiments using our algorithm as a post-processing unit in conjunction with the architecture, proposed in [51].
4.3.1 Quantitative Results
Using images from datasets provided by [35, 37] and [29], we compare the PSNR and SSIM metrics of our deblurred results with those of other state-of-the-art algorithms based on transformer architectures to determine if our prior enhances the efficacy of this class of algorithms.
Model Name RealBlur-J RealBlur-R HIDE GoPro Average
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
[51] 28.96 0.879 36.19 0.957 31.22 0.942 32.92 0.961 32.3225 0.93475
[52] 28.7 0.873 35.99 0.952 30.96 0.939 32.66 0.959 32.0775 0.93075
[41] 28.82 0.876 36.08 0.954 31.03 0.939 33.08 0.962 32.2525 0.93275
[5] 29.34 0.882 35.01 0.928 19.38 0.62 26.76 0.898 27.6225 0.832
[18] 27.97 0.834 33.79 0.903 24.51 0.871 28.7 0.858 28.7425 0.8665
[19] 28.7 0.866 35.26 0.944 26.61 0.875 29.55 0.934 30.03 0.90475
[55] 24.93 0.745 33.78 0.909 28.94 0.915 31.1 0.942 29.6875 0.87775
[53] 28.42 0.86 35.70 0.948 29.09 0.924 31.2 0.94 29.57 0.918
Ours 28.99 0.88 36.192 0.957 31.22 0.942 32.92 0.961 32.3305 0.935
Table 3: PSNR and SSIM scores for different models across various datasets
As evidenced by the data in Table 3, our algorithm, when employed as a post-processing unit, consistently delivers superior performance on average across all datasets and achieves the best results on two of these datasets individually
4.3.2 Qualitative Results
Beyond quantitative analysis, a critically important factor in assessing image quality is its visual appeal and ability to convince an audience of its superiority. To address this, we designed a study where we presented sets of images to a group of participants for comparison and scoring.
  1. 1.
    Type of Study: The study is an online survey. A base image was used to generate deblurred images using our model and models from [41, 52, 5, 53] and [51]. The deblurred images were concatenated in a random order in a 3x2 grid without labeling. We show grids containing 10 sets of images to our participants and ask them to score them out of 10. A higher score implies a better image.
  2. 2.
    Participants: The participants were chosen at random from a set of volunteer participants. The set of random participants belongs to no specific educational background.
  3. 3.
    Population Size: A total of 76 people participated in the survey.
  4. 4.
    Sampling Method: The performance of our model was tested against a number of datasets, from [35, 29, 37]. The sample set of 10 base images was picked at random with an equal distribution among all the datasets.
Taking the scores out of 10, here are the average ranks for the models included in the survey.
Figure 13: Comparison of our model with state-of-the-art models
Here, we see in Fig. 13 that the results from our model obtained the highest score (over 9 out of 10, on average) among all the models in the discussion.
5 Conclusions and Future Work
@@
In this paper, we propose a new prior for blind image deblurring, inspired by observations on the effects of convolution on image sparsity. We utilize Fourier transforms to extract frequency components, facilitating the separation of negative phases. This approach enabled the implicit determination of the blur kernel, informed by data on blur level and direction. The ReLU Sparsity prior is employed to penalize a higher L0 norm, which serves as an indicator of the blurriness of an image. To address the challenges associated with the non-convex optimization problem, we utilize half-quadratic splitting strategies. Our algorithm provides competitive performance on qualitative metrics such as PSNR, SSIM, and error ratio while providing upto three times lower inference times than the established state-of-the-art blind deblurring algorithms. Additionally, Our algorithm, serving as a post-processing unit, consistently outperforms competing methods, demonstrating its robustness and adaptability across a range of imaging conditions.
@@
For future work, we plan to explore the integration of our algorithm’s rapid inference and portable framework with deep learning architectures to further enhance deblurring evaluations. Additionally, we aim to develop alternative mathematical strategies to reduce the computational demands associated with gradient descent in RFT image computation. Another area of focus will be leveraging inherent image attributes such as saturation, intensity, and illumination to optimize the estimation of multiple kernels, thereby improving the recovery of the latent image without compromising processing speed.
Declaration
  • Funding: The authors did not receive support from any organization for the submitted work.
  • Competing interests: The authors have no competing interests to declare that are relevant to the content of this article.

### v3

@@ 13       $scale\leftarrow scale+1$
This blind deconvolution algorithm estimates the latent image $I$ and the blur kernel $k$, which is necessary for obtaining the latent image. Initially, the kernel $k$ is initialized to the coarsest resolution (Algorithm 1). Algorithm 1 determines the scale according to the current kernel size. The algorithm iteratively increases the scale. For each scale within a loop, the blurry image $B$ is down-sampled to the current scale in Algorithm 1. We iterate a predetermined number of times within this setting to refine the estimation of $I$ using Equation (6) in Algorithm 1, and $k$ using Equation (5) in Algorithm 1. Isolated noises are removed by deleting connected components within the kernel that do not exceed a specified threshold (Algorithm 1). The kernel is centered after the noise removal in Algorithm 1. the kernel is then up-sampled with the current scale in Algorithm 1. All down-sampling and up-sampling are done using bi-linear interpolation.
@@ 9 $I\_result\leftarrow I_{1}-$ bilateral_filter($\mathrm{diff}$)
For non-blind deconvolution, methods based on Laplacian priors [19] tend to be effective in preserving fine spatial details. We estimate an image $I_{1}$ with such priors in Algorithm 2 and Algorithm 2. Then, we use the L0 norm of the gradient from our equation Eq. 4 to estimate another latent image $I_{2}$ in Algorithm 2, since gradient-based methods have been shown to provide accurate results in suppressing ringing artifacts. Similar to ringing suppression methods discussed in [35], we compute a difference map between these two estimated images in Algorithm 2. We use bilateral filtering on the computed difference map and subtract the result from $I_{1}$ in Algorithm 2, thereby smoothing out artifacts and completing non-blind deconvolution.
4 Experiment
In this section, we discuss the parameter settings, quantitative results on datasets [18, 22], and qualitative results on images from different distributions.
4.1 Empirical Setting
@@
The parameters $\lambda,\mu$ and $\alpha$ are assigned values $0.0043\times 10^{-4},$ 0.004 and $2$, respectively. We compared the convergence of the algorithm for various values of these parameters with respect to kernel similarity (defined as SSIM between the estimated kernel and ground-truth kernel), and a more detailed analysis of this is provided in the supplementary material. The value of $max\_iter$ is set to be 5 as a trade-off between timing and precision. The non-blind deconvolution techniques mentioned by Pan et al. [32] are followed. We provide a Python codebase for blind image deblurring on single images. In order to ensure fairness in comparisons with pre-existing algorithms that were not implemented using Python, we avoided using GPU acceleration. However, since the previous works on blind image deblurring were implemented primarily on MATLAB, the translation to Python offers the opportunity for GPU access in the future as well as extension with deep learning methodologies.
4.2 Quantitative Results
@@
In this section, we evaluate our proposed algorithm with the state-of-the-art algorithms based on the metrics of PSNR, SSIM [40], and error ratio. Datasets from Köhler et al. [18] and Levin et al. [22] are widely-used benchmarks, and are used for evaluating our work against other blind image deblurring algorithms.
(a) PSNR comparison with other algorithms
(b) PSNR with vs without our prior
Figure 4: PSNR comparison for Köhler et al. dataset
@@
PSNR on Köhler et al. dataset: There are a total of 32 images. For each of the 4 images and 8 kernels, there are 199 corresponding ground truth images for computing PSNR. We calculated the results of our algorithm for all of these images and plotted the average for each of the 4 images and also for the average of the 4 images against other statistical deblurring algorithms from [35, 10, 20, 48, 33, 9, 43, 12] in Fig. 4(a). We also presented a comparison of PSNR results when our prior is used versus when it is not used in Fig. 4(b).
@@
PSNR SSIM Error ratio Average inference time (seconds)
[33] 27.54 0.8626 1.2076 109.6088
[34] 28.38 0.925 0.8776 15.0949
[42] 26.1235 0.83637 1.48 18.61
[4] 26.48693 0.8515 1.01234 65.2
Ours 28.34 0.887 1.78 8.057
@@
In Tab. 1, we show that our method achieves the best inference time while performing competitively in terms of PSNR, SSIM, and inference time.
Inference Time Comparison: We compare our work with [33] and [34] in this section, using datasets from Levin et al. [22], Köhler et al. [18] and Sun et al. [38], respectively. We report the inference times along with the image dimensions in the table below.
@@
$255\times 255$ $800\times 800$ $1024\times 800$
Pan et al. [33] 109.609 1550.932 2280.716
Pan et al. [34] 15.095 210.981 269.872
Wen et al. [42] 18.61 54.416 66.134
Chen et al. [4] 65.2 755.43 794.186
Ours 8.057 39.5764 55.72
Table 2: Inference Times of Algorithms on Images of Different Dimensions
@@
Our approach achieves significant speed gains: nearly two times lower compared to [34], as presented in Tab. 2.
4.3 Visual Results
In this section, we present the results of our algorithm on blurry images from different domains, which we compare with other state-of-the-art blind image deblurring algorithms. The images are taken from datasets provided by Lai et al. [21], Levin et al. [22] and Köhler et al. [18].
@@ 4.3 Visual Results
Figure 5: Results of our blind image deblurring algorithm, compared with other state-of-the-art algorithms. From left to right: 1) Input blurry image, followed by results from 2) Chen et al. [4] 3) Wen et al. [42] 4) Pan et al. [32], 5) Pan et al. [33] and 6) our algorithm.
@@
Figure 6: More results of our blind image deblurring algorithm, compared with other state-of-the-art algorithms. From left to right: 1) Input blurry image, followed by results from 2) Chen et al. [4] 3) Wen et al. [42] 4) Pan et al. [32], 5) Pan et al. [33] and 6) our algorithm.
We observe that our algorithm provides competitive results in terms of visual appeal, as well as qualitative measures mentioned in the previous section, for images irrespective of the distribution or domain they are taken from.
5 Conclusions and Remarks
@@
In this paper, we propose a new prior for blind image deblurring based on observations made on the effects of convolution on the sparsity of an image. We made use of Fourier transforms in order to obtain the frequency components in order to separate the negative phase(s), which enabled us to obtain the blur kernel implicitly through information on blur level and blur direction. The ReLU Sparsity prior can be used to penalize a higher L0 norm, which is an indicator of the blurriness of an image. For the non-convex optimization problem, we use half-quadratic splitting strategies. Our algorithm provides competitive performance on qualitative metrics such as PSNR, SSIM, and error ratio while providing up to two times lower inference times than the established state-of-the-art blind deblurring algorithms.
@@
Future work: We plan to explore the integration of our algorithm’s rapid inference and portable framework with deep learning architectures to further enhance deblurring evaluations. Additionally, we aim to develop alternative mathematical strategies to reduce the computational demands associated with gradient descent in RFT image computation. Another area of focus will be leveraging inherent image attributes such as saturation, intensity, and illumination to optimize the estimation of multiple kernels, thereby improving the recovery of the latent image without compromising processing speed.